### v1

# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 85 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP.  [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2024 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work.
Figure 1: Number of NLP papers using TDA published each year from 2013 to 2024
1 Introduction
Proliferation of the Internet has given rise to the generation of massive amounts of data. These massive amounts of data when processed can solve many crucial issues plaguing our current society. Due to this well-established notion among stake-holding institutions, the Machine Learning (ML) field has been thriving as a tool that extracts trends and solutions to non-trivial problems. However, real-world data tends to be noisy, heterogeneous, imbalanced, have missing labels, contain high-dimensionality, etc., often making the adoption of ML techniques to such datasets non-trivial. Therefore, to extract meaningful findings from data, specifically real-world data, clever techniques that extract additional features, while preserving the structure of the data need to be employed. To that end, a small niche community for Topological Data Analysis (TDA) applications in NLP has emerged. Being promised as a technique that can extract and analyze the shape/topology of data, TDA has great potential in mitigating such issues witnessed in real-world data.
TDA is a “collection of powerful tools that can quantify shape and structure in data” [^2] https://www.indicative.com/resource/topological-data-analysis/ and is inspired by the algebra and geometry mathematical fields. The benefits of TDA are vast, including the ability to extract additional features, typically not captured by other feature extraction techniques Uchendu et al. (2024); Michel et al. (2017). We call these features that TDA extracts - topological features. Unsurprisingly, since TDA is used to capture topological features, it has been applied to many tasks where data has distinct graphical structures Papamarkou et al. (2024); Hensel et al. (2021). These include tasks that have obvious graph-like structures such as protein classification Dey and Mandal (2018); Lamine et al. (2023); Valeriani et al. (2024) and drug discovery Alagappan et al. (2016); to those that are not so obvious such as diabetes classification Wamil et al. (2023); Skaf and Laubenbacher (2022), image classification Horn et al. (2022); Trofimov et al. (2023), time series analysis Petri and Leitao (2020); Tymochko et al. (2021); Gholizadeh and Zadrozny (2018), and speech processing Tulchinskii et al. (2023). However, since the shape of a text is not apparent, it has not gained as much attention in Natural Language Processing (NLP) as it has done in the Computer Vision Horn et al. (2022); Trofimov et al. (2023) and Medical Singh et al. (2023); Nielson et al. (2015) domains. Still, several researchers have found ways to extract additional features using TDA that the other typical numerical representation techniques in text such as TF-IDF, Word2Vec embeddings, BERT embeddings, etc have not captured. Figure 1 shows that there has been a gradual acceleration in the number of published works in TDA for NLP applications, and we project this trend will continue in the future.
TDA aims to answer the main question - what is the true shape of a data? We survey 85 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
2 Topological Data Analysis (TDA)
Topology is defined as “the study of geometric properties and spatial relations unaffected by the continuous change of shape or size of figures,” (Oxford Dictionary). TDA is then a collection of powerful techniques that can quantify the shape and structure of data Munch (2017). Two main techniques are used to extract TDA features - Persistent Homology and Mapper.
2.1 Persistent Homology
Figure 2: Illustration of the Persistent Homology technique using different radii to find the persistent features Rieck (2020). $\epsilon$ is the ball diameter.
Persistent Homology (PH) Edelsbrunner et al. (2008) is the most popular TDA technique. It uses algebraic topology methods to extract topological signatures at different spatial dimensions. This process involves representing data as a point cloud and performing deformation or perturbation processes to extract the true “shape” of data after the noise has been removed. To achieve this, PH employs Vietoris-Rips complex Munch (2017). Vietoris-Rips complex is a way to build simplicial complexes which are used to represent data in a topological space. A simplicial complex is a topological space built by putting points, lines, and higher dimensional shapes together. These formations reveal features that are holes in different dimensions, represented as betti numbers ($\beta_{d}$, $d$-dimension). Holes in the 0-dimension ($\beta_{0}$) is represented as one vertex, 1-dimension ($\beta_{1}$) is represented as an edge, 2-dimension ($\beta_{2}$) is represented as a triangle. Further, these features are called connected components, loops/tunnels, and voids, respectively.
Using the method described above, data is represented as a point cloud, and circles are drawn around each point. Next, the radius of each circle is increased using a defined range of points, such that if the circles get bigger and touch, one of the points disappears and this is recorded as a $death$. Additionally, this process of perturbation in different dimensions can cause the $birth$ of a new hole which is also recorded. Thus, due to these deformations, the following TDA features can be extracted and recorded in a 3-column matrix, which consists of columns representing - the $birth$ (formation of holes), $death$ (deformation or the closing of holes), and persistence features. Persistence is defined as the length of time it took a feature to disappear or die ($death-birth$). The $death$ is recorded with the radii value at which the points overlap. Lastly, TDA features are typically visualized in a persistence diagram which is a visual representation of the 3-column matrix of TDA features. Figure 2 illustrates an example of the process of extracting TDA features using Persistent Homology. Other ways of visualizing TDA features include Persistence images Adams et al. (2017) and Barcode plots Ghrist (2008). Application-wise, PH has been used to extract novel features to complement existing NLP representations and improve various text classification performances Doshi and Zadrozny (2018); Uchendu et al. (2024); Wu et al. (2022).
2.2 Mapper
Mapper is a dimension reduction clustering technique, used to visualize TDA-extracted topological structures/signatures. It was proposed by Singh et al. (2007) and has been used extensively to visualize topological structures in data to create visually pleasing figures, as well as interpret model performance through data probing Carlsson (2020). The Mapper algorithm works in 4 steps [^3] https://www.quantmetry.com/blog/topological-data-analysis-with-mapper/ (Figure 3) following Murugan and Robertson (2019)’s instructions: (1) Transform the data to a lower-dimensional space using a filter function $f$, also known as a lens. This implies projecting from one space to another. Options for filter functions include PCA Maćkiewicz and Ratajczak (1993), UMAP McInnes et al. (2018), and any other dimension-reduction algorithms; (2) Create a cover $(U_{i})_{i\in I}$ for the projected space, which is typically composed of overlapping intervals with a constant length; (3) Cluster the points in the preimage $f^{-1}(U_{i})$ into sets $C_{i,1},\ldots,C_{i,k_{i}}$ per interval $U_{i}$; (4) Create a graph where each vertex represents a cluster set, and there is an edge between two vertices if the corresponding clusters share common points. In this step, points in the same neighborhood are clustered using a defined clustering technique such as DBSCAN Ester et al. (1996) to change a cluster of several points into a node of a graph.
The intrinsic nature of the Mapper algorithm makes it advantageous in preserving structure even with mapping from one dimension to another. Furthermore, the clustering techniques allow it to be used to explain model performance as the clusters, and colors have meaning that can be further explored. In this survey, we will discuss how several researchers used it to explain several phenomena in NLP tasks.
Finally, Mapper is more useful for exploratory data analysis, while PH is more useful for analyzing point clouds and examining the persistence of features.
Figure 3: Illustration of Mapper from Murugan and Robertson (2019). The filter function $f$ is a height function, which is a projection onto the y-axis. The cover of the projected space is the four intervals $U_{i}$. The Mapper graph on the right is a result of applying the rest of the Mapper algorithm and clustering each preimage in the nearest neighbor.
3 Theoretical Approaches of TDA in NLP
Since the field of NLP is very interested in representing and analyzing texts or speech in meaningful ways, several theoretical approaches have been proposed to investigate how well these approaches align with linguistic principles. Thus to explain or confirm linguistic phenomena within the NLP paradigm, a few researchers have proposed topological approaches for probing NLP techniques. By employing TDA techniques - Persistent homology or Mapper to probe for linguistic phenomena, researchers aim to capture the topological space (both semantic and syntactic relationships) of texts, analyze and visualize topological topic evolution within texts, and extract the topological shape of words. See Table 1 in the Appendix for the theoretical approaches.
3.1 Topological Space
3.1.1 Semantic Topological Space
A semantic topological space is a conceptual framework used to represent and analyze the relationships between meanings (semantics) of words, phrases, or other linguistic units in a topological or shape structure. This representation involves mapping these units into a mathematical space where the distance or structure between them reflects semantic similarity or other relationships (i.e., Euclidean space $\to$ Topological space).
Due to the limitations of traditional clustering algorithms in capturing semantic relationships within texts, Chiang (2007) utilizes TDA and represents semantic space as a simplicial complex of Euclidean space. This technique captures the coherent topics of a document in the connected component dimension ($\beta_{0}$), representing the semantic topological space Chiang (2007). Similarly, Karlgren et al. (2014) visualizes the topological semantic space of text using Mapper which identifies the topical density of the space. To capture topological properties, they train two semantic spaces in a specific topical domain Karlgren et al. (2014). One space was trained only on articles of similar topics, and the other on introductory paragraphs of those same articles. Findings reveal that clusters of main concepts remained close for the space trained only on articles of similar topics. For the other topological space, the main concepts were randomly distributed Karlgren et al. (2014).
In addition, Cavaliere et al. (2017) extracts main concepts from the texts by probing the context-aware semantic topological space built with simplicial complexes. Finally, Wagner et al. (2012) uses TF-IDF to numerically represent the top 10-50 words in a corpus and build a topological space that analyzes the structure of similarities within several documents. This topological space is built using discrete Morse theory and persistent homology to find meaningful topological patterns Wagner et al. (2012). However, as of 2012, they found that their technique was unsuccessful due to the computational costs, which is a testament to how the NLP field has improved such that we now have more tractable solutions, such as dimensionality reduction algorithms McInnes et al. (2018), and TDA packages (Ripser Bauer (2021), Sklearn-TDA Saul and Tralie (2019), PHAT Bauer et al. (2017)), as well as compute resources to efficiently construct topological spaces from large or complex data.
3.1.2 Syntactic Topological Space
A syntactic topological space is a theoretical framework used to represent and analyze the relationships between syntactic structures from a topological perspective. This concept is particularly relevant in linguistics, where it helps model and understand the structural aspects of language such as grammar, sentence construction, or the hierarchical organization of those linguistic units.
Port et al. (2018) analyzes how syntactic parameters are distributed over different language families, including Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. Port et al. (2018) show that the three persistent connected components ($\beta_{0}$) in the Niger-Congo family represents its three subfamilies - Mande, Atlantic-Congo, and Kordofania. The syntactic topological structures of these languages also reveal the historical linguistic phenomena that the Hellenic branch played a role in the historical development of the Indo-European languages Port et al. (2018).
Similarly, Port et al. (2022) probes the interpretability of the syntactic topological space. This is done by incorporating more explanation of linguistic structures by introducing so-called homoplasy phenomena to explain persistent loops further. Homoplasy phenomena in syntax are observed when dissimilar languages exhibit syntactic similarities Port et al. (2022). Findings reveal that the Indo-European family languages - Czech, Lithuanian, Middle Dutch, and Swiss German have the same homoplasy phenomena Port et al. (2022).
3.2 Topology of Topic Evolution
The topology of topic evolution refers to the study and representation of how topics, themes, or concepts develop and change over time within a given corpus of texts or discourse in a topological space/structure. This concept is particularly relevant in fields where understanding the temporal dynamics of topics can provide insights into trends, shifts in public opinion, or the development of scientific or cultural themes.
Sami and Farrahi (2017) utilizes TDA to visualize the relationship between words in a text block, words in a corpus, and text blocks in a corpus. This is done by visualizing both local context (i.e., each text block in a set of sentences), and global context (i.e., occurrence of extracted words in the corpus) features. These features are extracted by using circular topology to represent words. Sami and Farrahi (2017) also find that using the circular topology in 2D space, core words from the text block or corpus stay close to the center, and the explanatory words remain close to the circle’s periphery.
3.3 Topological “Shape” of Words
The topological “shape” of words is a conceptual framework in linguistics and cognitive science that explores the structural properties of words. This framework leverages ideas from topology to capture the true shape of words in a linguistically meaningful way. Thus, using topological methods such as TDA, the structural properties of words can be extracted and analyzed.
To compare the language phylogenies or evolutionary history of language in the Indo-European language family, Draganov and Skiena (2024) investigates the “shape” of words by using their embeddings as proxies. Using persistent homology, they construct language phylogenetic trees for over 81 Indo-European languages. Experiments reveal that: (1) the shape of the word embedding of a language carries historical and structural information, similar to Port et al. (2018, 2022)’s findings; (2) TDA methods can successfully capture aspects of the shape of language embeddings Draganov and Skiena (2024). Fitz (2022) introduces a novel terminology - word manifold, which is a simplicial complex, whose topological space captures grammatical structure expressed by the corpus. This is done by implementing a technique for generating topological structure directly from strings of words. Fitz (2022). Experiments reveal that the homotopy type of the word manifold is also influenced by linguistic structure Fitz (2022). Similarly, Fitz et al. (2024) measures the topological complexity (known as perforation) of hidden representation of Transformer-based models to understand their topological shapes.
Similar to Port et al. (2018, 2022), Dong (2024) extracts the topological shapes of languages. Specifically South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families using TDA. By using techniques like multiple correspondence analysis (MCA) for dimension reduction of the categorical-valued dataset and persistent homology, Dong (2024) visualizes each language in the selected families as a point cloud. This forms the topological shape of the South American languages, such that languages close together are more similar. By comparing the topological shapes of the languages, it is observed that there are major distinctions between the Jê-proper and the non-Jê-proper languages, as well as the northern and southern Quechuan languages Dong (2024).
Figure 4: Illustration of the TDA feature extraction pipeline in NLP with three steps: 1-extracting numerical representations, 2-reformatting for TDA’s inputs, and 3-extracting TDA features.
Figure 5: Taxonomy of Topological Data Analysis (TDA) for Natural Language Processing (NLP) Applications
4 Non-theoretical Approaches of TDA in NLP
There are several ways to categorize the applied/non-theoretical TDA applications in NLP. These applications can be categorized by task, learning type, modality, TDA technique, and numerical representation. We observe that categorizing these TDA applications by task and numerical representation is more meaningful than the other categories since those categories are binary and not very descriptive of the landscape.
Out of these dimensions, the numerical representation showcases the bottleneck for extracting useful TDA features. See Figure 4 for an illustration of the pipeline for extracting TDA features from numerically represented texts. Therefore, while we focus on both task and numerical representation, our main taxonomy for the non-theoretical applications is centered on how TDA features are extracted from different forms of numerical representations. Figure 5 illustrates the taxonomy of non-theoretical applications of TDA in NLP tasks. See Table 2 in the Appendix for the non-theoretical approaches.
4.1 Tasks, Learning Types, Modality, and TDA techniques
We categorize the NLP tasks to which TDA has been applied into 7 categories: (1) Classification Zhu (2013); Tymochko et al. (2021); Tulchinskii et al. (2024); (2) Clustering and Topic Modeling Holmes (2020); Guan et al. (2016); (3) Sentiment and Semantic Analysis Michel et al. (2017); Rawson et al. (2022); Jakubowski et al. (2020); (4) Structure and Visualization Haghighatkhah et al. (2022); Rathore et al. (2023); (5) Health, Social, and Scholarly Analysis Effah (2017); Novak (2019); Petri and Leitao (2020); (6) Speech and Music Processing Bonafos et al. (2023); Sassone et al. (2022); and (7) Model Interpretation and Analysis Kostenok et al. (2023); Gourgoulias et al. (2024); Haim Meirom and Bobrowski (2022).
Recently TDA has been more heavily applied to five sub-tasks - deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024), linguistic/grammatical acceptability classification Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024); Perez and Reinauer (2022), word sense induction and disambiguation Rawson et al. (2022); Jakubowski et al. (2020); Tfemčinas (2018), and polysemy word classification Jakubowski et al. (2020); Triki (2021); Shehu (2024); Garcia (2022). All of these five subtasks are particularly interesting because the unique labels have very similar features, making them non-trivial to distinguish from each other. Therefore, researchers employed TDA as it is beneficial for extracting the true structure of a dataset, which could be useful for accurate classification. Lastly, we find that TDA is successful in such applications, achieving comparable performance to the baseline or mostly outperforming baselines.
Finally, we can also categorize non-theoretical TDA applications in NLP by learning types, Modality, and TDA techniques. Learning types has supervised Elyasi and Moghadam (2019); Lavery et al. (2024), and unsupervised Spannaus et al. (2024); Bonafos et al. (2024); Modality has text Triki (2021); Kostenok et al. (2023), and Speech Sassone et al. (2022); and TDA techniques, has Persistent Homology Torres-Tramón et al. (2015); Cherniavskii et al. (2022), and Mapper Holmes (2020); Elyasi and Moghadam (2019). Due to the binary nature of these three categories, they are not very informative and thus unsuitable as the main taxonomy for the non-theoretical applications.
4.2 Numerical Representation
4.2.1 TF-IDF
TF-IDF (Term Frequency - Inverse Document Frequency) is a well-known statistical formula that calculates the importance of words relative to a corpus. A few works investigated the extraction of topological features from TF-IDF representations as part of the pipeline illustrated in Fig. 4. For instance, SIFT, a persistent homology-based model with TF-IDF, is developed to differentiate between child and adolescent writings Zhu (2013). It is later adopted for deepfake text detection Løvlie (2023), extracting features from presidential election speeches from 2000-2020 Huang (2022), and movie genre classification with over 15% increase in F1 score Doshi and Zadrozny (2018); Shin (2019). Some other applications include - legal entailment Savle et al. (2019), sentiment analysis Michel et al. (2017), keyphrase extraction Guan et al. (2016), twitter topic detection Torres-Tramón et al. (2015), and text summarization Kumar and Sarkar (2022). In addition, TF-IDF has also been used as input for mapper to explain linguistic phenomena such as geographical differences of nursery rhymes Maadarani and Hajra (2020) and behaviors of NLP model van Veen (2020). Elyasi and Moghadam (2019) compares the two popular TDA approaches - Persistent Homology and Mapper to classify Persian poems. Lastly, Effah (2017) applies TDA to age group categorization of lonely people.
4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec Embeddings.
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Word2vec representations have been used as inputs to persistent homology to extract a hierarchical document structure, known as story trees, that allows tracing storylines Haghighatkhah et al. (2022). Word2vec has also been used with mapper to construct a graph of word2vec embeddings to create a topological search engine Cornell (2020). Next, Paluzo Hidalgo et al. (2019) employs TDA with word2vec to measure the distance between the literary style of Spanish poets belonging to similar genres - Francisco de Quevedo, Luis de Góngora, and Lope de Vega. TDA has also been applied on word2vec to solve innovative problems such as the detection of fraudulent papers Tymochko et al. (2021), detecting topological loops in logical statements Tymochko et al. (2020), and disease prediction from epidemic curves Petri and Leitao (2020).
In addition, Wu et al. (2022) using the TDA features extracted from Word2Vec representations, built an ensemble of BERT & TDA to detect contradiction types - Negation, Antonym, Replacement, Switch, Scope, and Latent. Yessenbayev and Kozhirbayev (2022) and Yessenbayev and Kozhirbayev (2024) compare embeddings extracted from text and speech using persistent homology. Holmes (2020), uses TDA to group embeddings for document clustering and topic modeling tasks. Rawson et al. (2022) and Tfemčinas (2018), build an unsupervised topological algorithm for word sense induction and disambiguation. Lastly, Feng et al. (2024) compares the data generated by traditional data augmentation techniques - Word2Vec and GloVe embeddings with LLM-generated augmented data (from GPT-J) using persistent homology.
GloVe Embeddings.
GloVe or Global Vectors for Word Representation is another technique for numerically representing texts as embeddings. GloVe has been used to extract useful topological features for classification (i.e., topological signatures of whole texts) such as those of novel writers Gholizadeh et al. (2018), time-series-based sliding windows Gholizadeh et al. (2020), and keywords Novak (2019). These features have also been used for unique applications such as comparing cross-lingual sentence representations Haim Meirom and Bobrowski (2022), explaining model performance Spannaus et al. (2024), social anxiety detection Byers (2021), and deepfake text detection Løvlie (2023). TDA features extracted from GloVe have also been compared with classical text embedding such as TF-IDF Michel et al. (2017) and also combined with those extracted from BERT and BiLSTM Deng and Duzhin (2022) in improving the performance of both classification and clustering tasks. Other applications include - extracting logic and reasoning in texts to test manifestation of intelligence and understanding in models Zadrozny (2021a), and capturing circles in circular arguments Tymochko et al. (2020); Zadrozny (2021b).
FastText Embeddings.
FastText embeddings build upon the Word2Vec approach by incorporating subword information, enhancing the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). Jakubowski et al. (2020), Triki (2021), and Shehu (2024) use persistent homology to measure the polysemy of a word. They find that topological polysemy is inconsistent across different word embeddings. Jakubowski et al. (2020) further uses a similar technique to perform word sense induction & disambiguation from SemEval-2010 task. Also, Tfemčinas (2018) performs unsupervised word sense induction & disambiguation. Tymochko et al. (2021) extracts topological features from FastText embeddings to detect fraudulent papers.
4.2.3 Pre-trained Contextual Embeddings
Transformer Embeddings.
Researchers evaluated the strength of the TDA features extracted from Transformer-based Vaswani et al. (2017) embeddings. Using the idea of self-attention, the neural network can encode more semantic and syntactic features than previous embeddings which should allow for richer TDA features to be extracted. To incorporate TDA features for classification, several researchers have investigated the efficacy of using other outputs of the encoder models - CLS Embedding output, hidden weights and attention weights to extract high-quality additional features.
CLS Embedding Output. Tulchinskii et al. (2024) proposes a persistent homology dimension estimator (PHD) that measures the intrinsic dimensionality of the manifold underlying the set of embeddings for a given text sample for deepfake text detection. Kushnareva et al. (2024) uses PHD to perform boundary detection (i.e., where the human author stops and the AI author begins) in the context of deepfake text detection. Lavery et al. (2024) performs fake news detection with a topologically-augmented BERT model. Proskura and Zaytsev (2024a) selects the best models to create an ensemble model with by first checking the topological similarity of the embeddings of the models. Das et al. (2021) extracts topological features using the sentence embeddings of BERT to represent texts as persistent image vectors to perform TEDtalk public speaking ratings classification. Gourgoulias et al. (2024) proposes an unsupervised approach for estimating the class separability of text datasets for binary and multi-class tasks. Rathore et al. (2023) proposes TopoBERT, a visual tool that topologically analyzes the fine-tuning process of the Transformer-based model using Mapper. Similarly, Byrne et al. (2022) uses Mapper for unsupervised topic modeling by representing the embeddings as a graph.
Hidden Weights. Uchendu et al. (2024) proposes TopFormer, an end-to-end fine-tuned Topological RoBERTa-base model for deepfake text attribution by using the reshaped $pooled\_output$ (to 2D) as input for the TDA layer. Similar to Jakubowski et al. (2020), Triki (2021), and Shehu (2024), Garcia (2022) studies polysemous words. Using Mapper to visualize the hidden representations of polysemous words in the BERT transformer model, Garcia (2022) studies how BERT represents the semantics of ambiguous words. Similarly, Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to describe the topological properties of internal representations in LLMs, detect topologically similar layers in an LLM, and prune redundant layers. Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT, correlates with test accuracy, and is robust to adversarial perturbations. García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification to encourage clustering of similar points. In addition, Ruppik et al. (2024) extracts topological features using persistent homology from RoBERTa’s hidden weights for dialogue term extraction. Finally, Kudriashov et al. (2024) probes BERT’s hidden weights using persistent homology when training BERT on new grammatical features, known as polypersonality.
Attention Weights. Attention weights of transformers have also been utilized as input features to extract TDA features. The majority of the applications focus on two models: BERT and RoBERTa. Regarding BERT model, its attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted that are useful for deepfake text detection Kushnareva et al. (2021) or investigation of whether language models learn grammatical concepts central to human linguistic competence (i.e., grammatical acceptability judgment) in different languages including English, Italian, Swedish, Russian Cherniavskii et al. (2022); Proskurina et al. (2023) and speech recognition task Tulchinskii et al. (2023). Similar techniques have also been used on RoBERTa’s attention weights for similar and other important tasks in linguistics such as grammatical checks Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), deepfake text detection Uchendu et al. (2024), and document coherence Jain et al. (2024).
We also notice a few works utilize the attention weights for analyzing different properties of NLP models such as model uncertainty estimation Kostenok et al. (2023), robustness evaluation of TDA features Perez and Reinauer (2022) and out-of-distribution detection (OOD) Pollano et al. (2024), dynamic weighting for ensemble models Proskura and Zaytsev (2024b). Especially,  Pollano et al. (2024) find that models utilizing TDA features extracted from attention weights perform best on far out-of-domain samples compared to when utilizing TDA features extracted from CLS outputs Pollano et al. (2024). This demonstrates that the choice of numerical representation used to extract TDA features is crucial.
ELMo Embeddings. ELMo embeddings (Embeddings from Language Models) are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) uses ELMo embeddings to extract topological features for detecting fraudulent papers by examining their titles and abstracts. To compare the strength of ELMo embeddings, they further extract topological features with other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
4.2.4 Symbolic Representations
Symbolic representations in the context of AI and cognitive science refer to the use of symbols such as letters, numbers, tokens, or abstract entities to represent concepts, objects, relationships, and rules within a system. These symbols can be manipulated according to predefined rules to perform reasoning, problem-solving, and decision-making. Symbolic representation contrasts with sub-symbolic representations, such as neural network-based embeddings, which do not explicitly use symbols or rules. This section then discusses the creation of symbolic representations by using principles of letter coding (PLC) and principles of speech sound coding (PSSC) which are then used to extract topological features.
PLC refers to rules and methods used to encode letters, which can be part of various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters. For instance, Yurchuk and Gurnik (2023) uses the PLC to create word embeddings for Ukrainian tongue twisters and extract topological features from such embeddings with persistent homology. This is to distinguish tongue twister from a simple narrative sentence using SVM and decision tree classifiers. Similarly, Kovaliuk et al. (2024) use PSSC for classifying spoken Ukrainian tongue twisters.
4.2.5 Multi-Modal Representations
TDA features have also been extracted from other NLP-related feature representations, including multi-media data such as audio and video. There have been a variety of research utilizing such multi-modal inputs to TDA layers, ranging from studying human vowels and infant vocalizations Bonafos et al. (2023, 2024); Noakes (1991), emotion recognition from audio speech Gonzalez-Diaz et al. (2019) and videos Paluzo-Hidalgo et al. (2022), depression detection from audio clips Tlachac et al. (2020) to recognizing voiced and voiceless consonants in speech Zhu et al. (2024) and music classification Bergomi (2015); Sassone et al. (2022), the majority of which utilizes persistent homology to extract TDA features. Due to the multi-modality nature of these works, they also help inform novel linguistic-sound findings such as consonants can be detected using vibration patterns that are captured by persistent homology Zhu et al. (2024).
5 Open Problems and Future Directions
5.1 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigates the shape of words and their embeddings in Indo-European languages and finds similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts, such that the languages clustered closely together are similar or one influenced the other. Finally, we observe only 11 theoretical TDA approaches in NLP, compared to the 70+ non-theoretical approaches. Thus, we need more theoretical TDA approaches as it is impractical to achieve the depth and understanding of performance from a topological perspective without more investigations.
5.2 TDA for Interpretability
Interpreting TDA features in NLP problems given its non-intuitive nature is very challenging. This is evident in the fact that most TDA for explainability applications is mostly in Computer Vision Saul and Arendt (2018), where the structure is distinct. Consequently, the interpretation of TDA features for text or speech data remains an open problem. There are currently two main tasks in this space - (1) explain model performance by interpreting TDA features; (2) explain model performance by using TDA to probe the prediction space or data. Either tasks require a deeper understanding of TDA such that intuitive explanations can be used to tie topology to linguistic phenomena. Specifically, we need novel approaches that link TDA features to linguistic phenomena, for instance, disentangling $\beta_{0}$, and $\beta_{1}$’s representations to different properties of natural texts such as coherency, and writing style. This can be done either through visualizing the prediction space of a model Rathore et al. (2023) or probing the prediction space of models with TDA Gardinazzi et al. (2024); Xenopoulos et al. (2022); Solunke et al. (2024).
5.3 Novel Applications of TDA
Once more theoretical approaches of TDA are investigated, and issues barring the application of TDA on interpretable NLP tasks are mitigated, then we hope that TDA can be applied to even more novel, diverse and important tasks. From Section 4.1, we can see that TDA has been applied to 7 categories on non-theoretical NLP tasks. While many of the tasks are interesting, especially the speech & music processing and health applications, there are still nuanced niche fields that could benefit from TDA. One glaring application is on multi-lingual tasks, just as Haim Meirom and Bobrowski (2022) and García-Castellanos et al. (2024). Due to the benefits of TDA which include performing robustly on heterogeneous, imbalanced, and noisy data, its application on multi-lingual tasks is necessary. Other applications include: semantic and syntactic structure, forensic authorship, embedding space, and multi-modal (e.g., language-vision model) analysis
5.4 Improvement in TDA Feature Extraction
Unlike, some other data modalities that have a distinct shape, texts take the shape of their numerical representation. This means that the shape of texts represented using TF-IDF will embody the TF-IDF representation, which tends to be different from other embeddings, such as the Transformer Pre-trained embeddings. While, the difference can be supported by the different linguistic features they capture such as semantic, syntactic, and lexical features, the differences in shape can be un-intuitive, making it difficult to determine the best numerical representation to extract the most useful TDA features. Therefore, we must find novel ways of representing texts numerically, for example, the use of symbolic representations or better ways to use the numerical representations that exist in such a way that is advantageous for extracting the best TDA features.
5.5 Adversarial Robustness of TDA Features
Robustness to noise, particularly adversarial perturbations, has been an important research topic in the NLP community. While it is encouraging that such robustness of TDA features is promising, there have been only a few works in this direction Perez and Reinauer (2022); Chauhan and Kaul (2022). For instance,  Perez and Reinauer (2022) show that their topologically-augmented BERT model is much more robust than the base BERT model when tested against perturbations generated by TextAttack Morris et al. (2020)Chauhan and Kaul (2022) also shows that there are some weak correlations between persistent homology features of a trained BERT model and its adversarial robustness against several state-of-the-art attack methods. Nevertheless, some results confirm that TDA is more robust, while others were inconclusive. Additionally, all existing works only evaluate on BERT model with simple attack mechanism, missing other security scenarios such as poisoned or backdoor attacks. Therefore, we call for a more comprehensive robustness evaluation of TDA on NLP models, especially from both NLP, ML and security community.
5.6 Topological Deep Learning for NLP
Due to the benefits of TDA and deep learning, a new niche field is born - Topological Deep Learning (TDL). Initially, TDL is described as an ensemble of topological features extracted by TDA techniques such as persistent homology and deep learning features. However, as the field has advanced, a new definition for TDL has emerged - “the collection of ideas and methods related to the use of topological concepts in deep learning” Papamarkou et al. (2024). This allows a deep learning model to be integrated more deeply with concepts of algebraic topology, such as the introduction of simplicial neural networks (NN), which are NNs with layers made up of simplicial complexes. This deeper integration of TDA into NNs makes TDL particularly useful for the explosion of data we are currently experiencing, especially because these datasets are now more often high-dimensional.
These high-dimensional data need better tools for processing as the current tools shrink the dimension, encouraging loss of information. In NLP, one particular approach to integrate TDA with high-dimensional NLP embeddings has been the utilization of text in graphical forms, which have been shown to yield better results than directly using texts as a sequence of tokens Zhong et al. (2020); Liu et al. (2023); Phan et al. (2023). More research is still needed to validate such an approach to a variety of NLP tasks and problems.
6 Conclusion
We surveyed 85 papers that applied TDA, using theoretical and non-theoretical approaches in NLP. In the theoretical approaches, researchers are more concerned with explaining linguistic phenomena from a topological perspective. For the non-theoretical approaches, several categories can be used for the taxonomy. Some are binary categories, yielding uninformative distinctions on the non-theoretical applications of TDA in NLP, these include - TDA techniques which are Persistent Homology and Mapper; data modality which is text and speech; and learning types which are supervised and unsupervised. While the more informative categories are tasks and numerical representation. Tasks are divided into 7 categories. Next, the main taxonomy for the non-theoretical approaches is the numerical representation used on texts before extracting TDA features. This is because, in NLP the bottleneck for extracting useful TDA features is the quality of the numerical representation techniques. Finally, based on the utility of TDA in NLP, we present the first survey which discusses the novel applications of TDA in NLP, as well as the challenges and open problems observed.
Acknowledgments
Adaku would like to thank Dr. Charlie Dagli for his encouragement, reading the paper drafts, and providing invaluable recommendations.
Appendix A Appendix
Name Category Task TDA Technique
Chiang (2007) Topological Space (Semantic) Capturing semantic relationships within texts Persistent Homology
Karlgren et al. (2014) Topological Space (Semantic) Identify topical density of the space Mapper
Cavaliere et al. (2017) Topological Space (Semantic) Extracting extracts main concepts from text Persistent Homology
Wagner et al. (2012) Topological Space (Semantic) Analyzes document similarities Persistent Homology
Port et al. (2018) Topological Space (Syntactic) Analyzes syntactic parameters of different language families Persistent Homology
Port et al. (2022) Topological Space (Syntactic) Explaining linguistic structures with homoplasy phenomena Persistent Homology
Sami and Farrahi (2017) Topology of Topic Evolution Topic evolution within documents Persistent Homology
Draganov and Skiena (2024) Topological “Shape” of Words Investigate the “shape” of language phylogenies in the Indo-European language family Persistent Homology
Fitz (2022) Topological “Shape” of Words Capturing grammatical structure expressed by corpus using word manifold Persistent Homology
Fitz et al. (2024) Topological “Shape” of Words Measure of topological complexity of Transformer-based hidden representations Persistent Homology
Dong (2024) Topological “Shape” of Words Investigates the topological shapes of South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families Persistent Homology
Table 1: Theoretical Applications of TDA in NLP.
Name Task Numerical Representation Learning Type Modality TDA Technique
SIFT Zhu (2013) child vs. adolescent writing detection TF-IDF Supervised Text Persistent Homology
Løvlie (2023) deepfake text detection TF-IDF, GloVe Supervised Text Persistent Homology
Huang (2022) election speech feature extraction TF-IDF Supervised Text Persistent Homology
Doshi and Zadrozny (2018); Shin (2019) movie genre classification TF-IDF Supervised Text Persistent Homology
Savle et al. (2019) legal document entailment TF-IDF Supervised Text Persistent Homology
Michel et al. (2017) clustering and sentiment analysis TF-IDF, GloVe Supervised Text Persistent Homology
DoCollapse Guan et al. (2016) keyphrase extraction TF-IDF Unsupervised Text Persistent Homology
TOPOL Torres-Tramón et al. (2015) Twitter topic detection TF-IDF Supervised Text Persistent Homology
Kumar and Sarkar (2022) text summarization TF-IDF Unsupervised Text Persistent Homology
van Veen (2020) Propaganda tweet classification TF-IDF Unsupervised Text Mapper
Elyasi and Moghadam (2019) classify Persian poems TF-IDF Supervised Text Persistent Homology & Mapper
Effah (2017) age group categorization of lonely people TF-IDF Supervised Text Persistent Homology & Mapper
Maadarani and Hajra (2020) nursery rhyme classification from different continents - Australia, Asia, Africa, Europe, and North America TF-IDF Supervised Text Persistent Homology
Haghighatkhah et al. (2022) building document structure Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
BERT+TDA Wu et al. (2022) contradiction detection Pre-trained (Word2Vec) Supervised Text Persistent Homology
Yessenbayev and Kozhirbayev (2022) speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech Persistent Homology
Yessenbayev and Kozhirbayev (2024) speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech Persistent Homology
Cornell (2020) building a topological search engine Pre-trained (Word2Vec) Unsupervised Text Mapper
Holmes (2020) document clustering and topic modeling tasks Pre-trained (Word2Vec) Supervised Text Mapper
Rawson et al. (2022) word sense induction and disambiguation Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
Tfemčinas (2018) word sense induction and disambiguation Pre-trained (Word2Vec, GloVe) Unsupervised Text Persistent Homology
Feng et al. (2024) Geometry of textual data augmentation Pre-trained (Word2Vec) Supervised Text Persistent Homology
Tymochko et al. (2021) fraudulent paper detection Pre-trained (Word2Vec, GloVe, ElMo) Supervised Text Persistent Homology
Petri and Leitao (2020) disease epidemic prediction Pre-trained (Word2Vec) Supervised Text Persistent Homology
Tymochko et al. (2020) finding topological loops in logical statements Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
Paluzo Hidalgo et al. (2019) measuring the distance between the literary style of Spanish poets Pre-trained (Word2Vec) Supervised Text Persistent Homology
Gholizadeh et al. (2018) extract the topological signatures of novelists Pre-trained (GloVe) Supervised Text Persistent Homology
TIES Gholizadeh et al. (2020) document categorization & sentiment analysis Pre-trained (GloVe) Unsupervised Text Persistent Homology
Spannaus et al. (2024) phenotype prediction and news group categorization Pre-trained (GloVe) Unsupervised Text Mapper
Zadrozny (2021b) finding topological loops in logical statements Pre-trained (GloVe) Unsupervised Text Persistent Homology
Byers (2021) social anxiety detection Pre-trained (GloVe) Supervised Text Persistent Homology
Haim Meirom and Bobrowski (2022) compare cross-lingual sentence representations Pre-trained (GloVe) Unsupervised Text Persistent Homology
Zadrozny (2021a) investigates the manifestations of intelligence and understanding in neural networks Pre-trained (Transformers - GloVe) Supervised Text Persistent Homology
Deng and Duzhin (2022) fake news detection Pre-trained (GloVe, BERT) Supervised Text Persistent Homology
Novak (2019) analyzing scholarly network Pre-trained (GloVe, BERT) Unsupervised Text Persistent Homology
Jakubowski et al. (2020) (1) polysemy word classification, and (2) word sense induction & disambiguation Pre-trained (FastText) Unsupervised Text Persistent Homology
Triki (2021); Shehu (2024) polysemy word classification Pre-trained (FastText) Supervised Text Persistent Homology
PHD Tulchinskii et al. (2024) deepfake text detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Kushnareva et al. (2024) deepfake text detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Gourgoulias et al. (2024) class separability estimation Pre-trained (Transformers - CLS) Unsupervised Text Persistent Homology
TopoBERT Rathore et al. (2023) visually analyzing the fine-tuning process of a Transformer-based model Pre-trained (Transformers - CLS) Unsupervised Text Mapper
Lavery et al. (2024) fake news detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Proskura and Zaytsev (2024a) Ensemble model selection Pre-trained (Transformers - CLS) Unsupervised Text Persistent Homology
Das et al. (2021) classification of public speaking ratings from TED talks Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Byrne et al. (2022) topic modeling Pre-trained (Transformers - CLS) Unsupervised Text Mapper
TOPFORMER Uchendu et al. (2024) deepfake text detection Pre-trained (Transformers - pooled_output) Supervised Text Persistent Homology
Garcia (2022) polysemy word classification Pre-trained (Transformers - Hidden) Supervised Text Mapper
Persistent Similarity Gardinazzi et al. (2024) Probing layers in LLMs Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Persistence Scoring Function Chauhan and Kaul (2022) captures the homology of the high-dimensional hidden representations Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Topological Densification García-Castellanos et al. (2024) zero-shot model stitching Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Ruppik et al. (2024) dialogue term extraction Pre-trained (Transformers - Hidden) Supervised Text Persistent Homology
Kudriashov et al. (2024) polypersonality Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Kushnareva et al. (2021) deepfake text detection Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024) grammatical acceptability judgment Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Kostenok et al. (2023) Uncertainty estimation of model predictions Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Perez and Reinauer (2022) spam detection, grammatical acceptability judgment, and movie sentiment analysis Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Pollano et al. (2024) out-of-distribution detection Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology
Proskura and Zaytsev (2024b) estimation of weights for ensembles of classification models Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology
TopoHuBERT Tulchinskii et al. (2023) speaker recognition Pre-trained (Transformers - Attention) Supervised Speech Persistent Homology
Vukovic et al. (2022) dialogue term extraction Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Yurchuk and Gurnik (2023) Ukrainian tongue twisters classification Symbolic Representations Supervised Text Persistent Homology
Kovaliuk et al. (2024) Ukrainian tongue twisters classification Symbolic Representations Supervised Speech Persistent Homology
Bonafos et al. (2023) human vowel classification Multi-Modal Supervised Speech Persistent Homology
Bonafos et al. (2024) infant vocalization classification Multi-Modal Unsupervised Speech Persistent Homology
Gonzalez-Diaz et al. (2019); Paluzo-Hidalgo et al. (2022) emotion recognition Multi-Modal Supervised Speech Persistent Homology
Tlachac et al. (2020) depression detection Multi-Modal Supervised Speech Persistent Homology
Zhu et al. (2024) consonants recognition Multi-Modal Supervised Speech Persistent Homology
Sassone et al. (2022) music genre classification Multi-Modal Supervised Speech Persistent Homology
Bergomi (2015) music classification Multi-Modal Supervised Speech Persistent Homology
Table 2: Non-theoretical applications of TDA in NLP